# Momentum-space Metric Tensor from Nonadiabatic Evolution of Bloch Electrons

### v1

@@ Abstract
A concrete example is the harmonic potential
$V(\bm{x})=\frac{\kappa}{2}\bm{x}^{2},$ (14)
which leads to $x^{i}=-\frac{1}{\kappa}\dot{q}^{i}$ and $\dot{x}^{i}=-\frac{1}{\kappa}\ddot{q}^{i}$. Here $x^{i}=\delta^{ij}x_{j}$ and $\delta^{ij}$ the Kronecker delta. The equation of $\bm{q}$ is that in Eq. (13) with $\bar{G}_{ij}=G_{ij}+\frac{1}{\kappa}$. This equation follows from the Lagrangian [59]
$L=\frac{1}{2}\bar{G}_{ij}\dot{q}^{i}\dot{q}^{j}-A_{i}\dot{q}^{i}-E(\bm{q}),$ (15)
where $E(\bm{q})$ is a scalar potential and $A_{j}$ is a vector potential. Unlike a relativistic particle in curved spacetime, where $\bar{G}_{ij}\dot{q}^{i}\dot{q}^{j}=\text{constant}$ due to the constant seed of light, no such constraint applies here.
@@
This Lagrangian has a well-defined canonical structure enabling us to perform canonical quantization. The canonical momentum is
$p_{i}=\frac{\partial L}{\partial\dot{q}^{i}}=\bar{G}_{ij}\dot{q}^{j}-A_{i}.$ (16)
In specific, with a constant nonadiabatic metric $G_{ij}=G_{0}\mathcal{I}_{ij}$ with $\mathcal{I}$ the identity matrix, an effective mass emerges and the Hamiltonian simplifies to
$H=\frac{1}{2M}\delta^{ij}(p_{i}+A_{i})(p_{j}+A_{j})+E(\bm{q})$ (17)
describing the dynamics of a massive particle under electromagnetic fields on a flat torus with $M=G_{0}+\frac{1}{\kappa}$. The covariant Laplacian will need to be used when $G_{ij}$ is not a constant.
@@
By promoting $(p_{i},q^{j})$ to operators with the commutator $[\hat{p}_{i},\hat{q}^{j}]=-i\delta_{i}^{j}$, we obtain the quantized Hamiltonian. The re-quantization of semiclassical wave packet dynamics is fruitful in describing various quantum phenomena such as Landau levels of Bloch electrons [6] and Berry curvature effects on exciton spectrum [47]. We note that here ${\bm{x}}=-\frac{1}{\kappa}\dot{\bm{q}}=-\frac{1}{1+\kappa G_{0}}({\bm{p}}+{\bm{A}})$, consistent with the canonical variables in Ref. 47, but with the distinction that it incorporates a nonadiabatic correction from $\kappa G_{0}$.
The topological structure of the Brillouin zone as a torus requires that the $\bm{q}$-space wave functions satisfy periodic boundary conditions. This contrasts with previous approaches [6, 47], which performed re-quantization in real-space coordinates without imposing such conditions. This boundary condition is compatible with the quantization of the Chern number, enabling a direct connection between the electronic states in $\bm{q}$-space to the quantum Hall effects on a torus, as discussed below.
In the presence of a time-dependent uniform electric field, the harmonic potential acquires a time-dependent center, leading to a time-dependent vector potential in the effective Hamiltonian that encodes the influence of the electric field.
This $\bm{q}$-space framework provides an alternative lens for studying Bloch electron dynamics under static or dynamic confining potentials, a class of problems that are particularly interesting in flat-band systems, where finite mass and geodesic motion arise from the interplay between the potential and the metric $G_{ij}$.
Single Particle Dynamics in a Flat Band— The scalar potential $E(\bm{q})$ is a constant in flat bands, and the electron dynamics are determined by the Berry curvature and nonadiabatic metric. Here we study a two-dimensional flat band. We focus on the simplest case when the Berry curvature $F_{ij}$ and nonadiabatic metric $G_{ij}$ are constants, and leave the non-uniform case for future study. This simplifies the dynamics from a curved manifold to a Euclidean torus. Consider a rectangular Brillouin zone with dimensions $L_{q_{x},q_{y}}=2\pi/L_{x,y}$ where $L_{q_{x},q_{y}}$ defines the ${\bm{q}}$-space size while $L_{x,y}$ are the real-space unit cell dimensions.
When the particle in this flat band is confined in a harmonic potential shown in Eq. (14), the Hamiltonian in $\bm{q}$-space follows Eq. (17)
$\hat{H}=\frac{1}{2M}(\hat{\bm{p}}+\hat{\bm{A}})^{2},$ (18)
where $M=G_{0}+\frac{1}{\kappa}$ for a constant metric $G_{ij}=G_{0}\mathcal{I}_{ij}$, and $\hat{\bm{A}}=\frac{1}{2}F\hat{z}\times\hat{\bm{q}}$, corresponding to $F_{12}=-F_{21}=-F$. This describes a Schrödinger equation with periodic boundary conditions in $\bm{q}$-space.
The Berry curvature’s integral over the Brillouin zone corresponds to the Chern number $C$, i.e., $FL_{q_{x}}L_{q_{y}}=2\pi C$. This indicates a quantized effective magnetic flux generated by $C$ Dirac monopoles within the torus. This ensures the self-consistency of the equation, i.e., the existence of solutions satisfying the boundary condition.
@@
With nonzero $C$, the eigenvalue problem of Eq. (18) becomes a Landau level problem on a torus [81, 82]. The eigenfunctions are Landau level wavefunctions on a torus. The eigenenergies are $\epsilon_{n}=(n+\frac{1}{2})\frac{F}{M}=(n+\frac{1}{2})\frac{F\kappa}{1+\kappa G_{0}}$ with each level’s degeneracy equal to $|C|$. The confining potential $\kappa$ corresponds to an effective inverse mass, whereas the nonadiabatic metric contributes to a correction from interband mixing. The interband mixing effect is enhanced by increasing the potential strength that increases $\dot{\bm{q}}$ and thus the interband contributions.
Alternative to TKNN’s approach, which employs a periodic potential to lift Landau level degeneracy [83], a harmonic potential can also lift the degeneracy of a flat band. This maps the flat-band particle within a confining potential in real space into a Landau level problem on a torus in $\bm{q}$-space, allowing for analytical solutions of both eigenfunctions and eigenenergies. The nonadiabatic metric induces corrections from interband mixing. The metric contribution is significant when $|\kappa G_{0}|\gg 1$.
@@
Two-body Bound States in Flat Bands— The effect of $G_{0}$ can be enhanced in a two-body problem [47] where $\kappa$ represents the Coulomb interaction strength. Here we consider two wave packets in flat bands. We assume that the wave packets are well separated, and they interact through a softened Coulomb interaction in real space. The two-particle Lagrangian is
@@
where $\hat{\bm{q}}_{-}=(\hat{\bm{q}}_{1}-\hat{\bm{q}}_{2})/\sqrt{2}$, $[\hat{p}_{-,i},\hat{q}_{-}^{j}]=-i\delta_{i}^{j}$, and $M^{\prime}=G_{0}+\frac{1}{2\kappa_{\rm I}}$ is the reduced effective mass. This Hamiltonian again resembles a Landau level problem on a torus (see Fig. 3) with eigenenergies $\epsilon_{n}=(n+\frac{1}{2})\frac{F}{M^{\prime}}-V_{0}$.
If the two particles are different, like an electron and a hole, the wave function of the relative coordinate is only restricted by the periodic boundary condition. If they are indistinguishable, the two-body wavefunctions need to be antisymmetric for fermions or symmetric for bosons. This leads to another boundary condition in addition to the periodic boundary conditions. Therefore, while the degeneracy of eigen-energy levels obtained by solving Eq. (21) is the Chern number $|C|$ ($C\neq 0$), only those states satisfying both boundary conditions are allowed solutions.
@@
Summary— We introduce a geometric framework that extends the study of Bloch electron dynamics beyond adiabatic Berry-phase effects. We show that the leading-order corrections to the adiabatic dynamics, the nonadiabatic effects, are elegantly characterized by the nonadiabatic metric—a momentum-space metric tensor in semiclassical wave packet dynamics. This metric unifies the nonlinear and nonadiabatic responses by introducing geometric and geodesic velocities beyond the anomalous velocity from the Berry curvature, enabling further generalization. Fundamentally, the nonadiabatic metric endows momentum space with a curved geometry, making the electronic dynamics a forced geodesic motion in momentum space. The momentum space picture provides an intuitive understanding of bound-state problems of flat-band particles, where massive dynamics is retrieved and the nonadiabatic metric captures interband mixing effects. Notably, the two-body bound-state wave functions in flat Chern bands resemble Landau level states on a torus.
Acknowledgements.
@@
Acknowledgments— Y.R. is indebted to Qian Niu, Yang Gao, Di Xiao, Wenqin Chen and Jie Wang for useful discussions.

### v2

@@ II.3 Effective Mass Corrections for Bound States
where $\bar{G}_{ij}=G_{ij}+\frac{1}{\kappa}$, $\bar{G}^{ij}\bar{G}_{jk}=\delta^{i}_{k}$, and $\bar{\Gamma}^{i}_{jk}$ is the corresponding Christoffel symbol that vanishes in this example as $G_{ij}$ is constant. $F_{jk}$ corresponds to the artificial magnetic field in the momentum space from Berry curvature. This equation follows the Lagrangian [58]
$L=\frac{1}{2}\bar{G}_{ij}\dot{q}^{i}\dot{q}^{j}-A_{i}\dot{q}^{i}$ (14)
@@
where $A_{j}$ is a vector potential and the band energy is a constant that is set as zero. This Lagrangian has a well-defined canonical structure enabling us to perform canonical quantization. The canonical momentum is $p_{i}=\frac{\partial L}{\partial\dot{q}^{i}}=\bar{G}_{ij}\dot{q}^{j}-A_{i}$ and the Hamiltonian reads
$H=\frac{1}{2M}\delta^{ij}(p_{i}+A_{i})(p_{j}+A_{j})$ (15)
@@
describing the dynamics of a massive particle under electromagnetic fields on a flat torus with $M=G_{0}+\frac{1}{\kappa}$. By promoting $(p_{i},q^{j})$ to operators with the commutator $\hat{p}_{i},\hat{q}^{j}]=-i\delta_{i}^{j}$, we obtain the quantized Hamiltonian in momentum-space representation. We note that here the canonical momentum $\hat{p}_{i}$ corresponds to the real-space coordinate, ${\bm{x}}=-\frac{1}{\kappa}\dot{\bm{q}}=-\frac{1}{1+\kappa G_{0}}({\bm{p}}+{\bm{A}})$, consistent with the canonical variables in Ref. 30, but with the distinction that it incorporates a nonadiabatic correction from $\kappa G_{0}$.
@@
The topological structure of the Brillouin zone as a torus requires that the $\bm{q}$-space wave functions satisfy periodic boundary conditions. This boundary condition is compatible with the quantization of the Chern number, enabling a direct connection between the electronic states in $\bm{q}$-space to the quantum Hall effects on a torus [84, 85]. The eigenenergies are $\epsilon_{n}=(n+\frac{1}{2})\frac{F}{M}=(n+\frac{1}{2})\frac{F\kappa}{1+\kappa G_{0}}$ with each level’s degeneracy equal to $|C|$. The confining potential $\kappa$ corresponds to an effective inverse mass, whereas the nonadiabatic metric contributes to a correction from interband mixing. The interband mixing effect is enhanced by increasing the potential strength that increases $\dot{\bm{q}}$ and thus the interband contributions.
@@
The confining problem can also be generalized to two-body bound states [30] where the effect of $G_{0}$ can be enhanced as $\kappa$ originates from the Coulomb interaction strength. Here we consider two wave packets in flat bands. We assume that the wave packets are well separated, and they interact through a softened Coulomb interaction in real space. The two-particle Lagrangian is
@@
where $\hat{\bm{q}}_{-}=(\hat{\bm{q}}_{1}-\hat{\bm{q}}_{2})/\sqrt{2}$, $\hat{p}_{-,i},\hat{q}_{-}^{j}]=-i\delta_{i}^{j}$, and $M^{\prime}=G_{0}+\frac{1}{2\kappa_{\rm I}}$ is the reduced effective mass. This Hamiltonian again resembles a Landau level problem on a torus (see Fig. 3) with eigenenergies $\epsilon_{n}=(n+\frac{1}{2})\frac{F}{M^{\prime}}-V_{0}$.
If the two particles are different, like an electron and a hole, the wave function of the relative coordinate is only restricted by the periodic boundary condition. If they are indistinguishable, the two-body wavefunctions need to be antisymmetric for fermions or symmetric for bosons. This leads to another boundary condition in addition to the periodic boundary conditions. Therefore, while the degeneracy of eigen-energy levels obtained by solving equation (18) is the Chern number $|C|$ ($C\neq 0$), only those states satisfying both boundary conditions are allowed solutions.
@@
II.4 Metric Tensor and Forced Geodesic Equation
The nonadiabatic metric also underscores the importance of the $\bm{q}$-space picture, instead of the real space, where $\bm{q}$ is taken as a generalized coordinate to describe the wave packet dynamics. This enables us to study the emergent gravity and emergent mass in momentum space, which will be discussed in this and following sections, separately.
$G_{ij}(\bm{q})$ in the Lagrangian defines a metric tensor in the tangent vector space of the generalized velocity $\dot{\bm{q}}$ defining a line element $ds^{2}=G_{ij}dq^{i}dq^{j}$. In contrast to the quantum metric, which defines the quantum distance $ds_{Q}$ between two wave functions $|u_{\bm{q}}\rangle$ and $|u_{\bm{q}+d\bm{q}}\rangle$ at $\bm{q}$ and $\bm{q}+d\bm{q}$ via $ds_{Q}^{2}=1-|\langle u_{\bm{q}}|u_{\bm{q}+d\bm{q}}\rangle|^{2}$ [3], the nonadiabatic metric defines the geometric distance between the momentum-space centers of two wave packets. A $\bm{q}$-dependent $G_{ij}$ signifies a curved manifold. Conversely, in cases where $G_{ij}$ is constant across the momentum space, it corresponds to an effective mass in a flat $\bm{q}$-space, simplifying the dynamics to a Euclidean framework.
where $\Gamma^{l}_{jk}$ is the Christoffel symbol of the second kind
$G^{ji}G_{ik}=\delta^{j}_{k}$ with $\delta^{j}_{k}$ the Kronecker delta. When the right-hand side of equation (19) vanishes, the equation reduces to the geodesic equation for a free particle navigating a curved manifold with metric tensor $G_{ij}$. A nonzero right-hand side indicates a particle subject to forces from three sources (see Fig. 3): a scalar potential from the band dispersion $E$, a gauge field from the Berry curvature $F_{ij}$, and an external force from the wave packet velocity $\dot{x}_{j}$. This equation enables the study of gravity in condensed matter setups [92]. It is noted that equation (19) differs from the recent study on momentum space gravity [93, 94, 95] where $G^{ij}$ and $\ddot{\bm{q}}$ are absent and $\Gamma_{i,jk}$ instead of $\Gamma^{i}_{jk}$ is used.
When $x_{i}$ can be invertibly represented as a function of $\dot{q}^{j}$ by using $\dot{q}^{i}=-\frac{\partial E}{\partial x_{i}}$, we have
the equations of motion find their $\bm{q}$-space representation:
III Discussions
@@
We introduce a geometric framework that extends the study of Bloch electron dynamics in a scalar potential beyond adiabatic Berry-phase effects. We show that the leading-order corrections to the adiabatic dynamics, the nonadiabatic effects, are elegantly characterized by the nonadiabatic metric—a momentum-space metric tensor in semiclassical wave packet dynamics, distinguishing them from the conventional quantum metric. This metric unifies the nonlinear and nonadiabatic responses by introducing geometric and geodesic velocities beyond the anomalous velocity from the Berry curvature, enabling further generalization. It also modifies the quantum electron dynamics, manifested as a metric in momentum space. A constant nonadiabatic metric influences the flat-band electron dynamics in a harmonic potential by introducing a correction to the effective mass for the generalized coordinate $\bm{q}$. Depending on the origin of the harmonic potential, this mass correction influences the electronic properties in static potentials, Wigner crystal [78, 81], or flat-band excitons [30]. In general, the nonadiabatic metric is not a constant, which endows momentum space with a curved geometry, making the electronic dynamics a forced geodesic motion in momentum space. This enables the study of momentum-space gravity.
IV Methods
IV.1 Integrate out $c_{n}$
Given the Lagrangian
$L$ $=\sum_{n}c_{n}^{*}i\dot{c}_{n}-\dot{\bm{q}}_{c}\cdot\bm{x}_{c}-E_{c}+\Lambda_{[00]}+\sum_{n\neq 0}-c_{n}^{*}c_{n}\left[\omega_{[n0]}-(\Lambda_{[nn]}-\Lambda_{[00]})\right]+(c_{0}^{*}c_{n}\Lambda_{[0n]}+c.c.)$
where $c_{0}$ is the reference-band amplitude and $c_{n}$ are dynamical variables similar to $\bm{x}_{c}$ and $\bm{q}_{c}$. The normalization implies $c_{0}=1+\mathcal{O}(c_{n\neq 0}^{2})$ and the phase of $c_{0}$ contributes a total derivative and can be gauged away. For simplicity, we define
$\Delta_{n}\equiv\omega_{[n0]}-(\Lambda_{[nn]}-\Lambda_{[00]}),\qquad J_{n}\equiv c_{0}\Lambda_{[n0]},~n\neq 0.$ (23)
The action $S=\int dtL$ is quadratic in the excited-band amplitudes $c_{n\neq 0}$. Varying S with respect to $c_{n}^{*}$ gives the saddle-point equation
$(i\partial_{t}-\Delta_{n})c_{n}=-J_{n}\qquad n\neq 0.$ (24)
[…]
$S_{\rm eff}[\bm{q}_{c},\bm{x}_{c}]$ $=\int dtL_{0}(\bm{q}_{c},\bm{x}_{c})$
$-\sum_{n\neq 0}\int dtdt^{\prime}J_{n}^{*}(t)G_{n}(t,t^{\prime})J_{n}(t^{\prime}),$ (25)
where $L_{0}(\bm{q}_{c},\bm{x}_{c})\equiv-\dot{\bm{q}}_{c}\cdot\bm{x}_{c}-E_{c,0}+\Lambda_{[00]}$.
When parameters $\bm{q}_{c}$ vary slowly, we expand the inverse operator:
$G_{n}$ $=(i\partial_{t}-\Delta_{n})^{-1}\delta(t-t^{\prime})$
$=\left[-\frac{1}{\Delta_{n}}+\frac{i}{\Delta_{n}^{2}}\partial_{t}+\mathcal{O}(\partial_{t}^{2})\right]\delta(t-t^{\prime}),$ (26)
and keep terms up to second order in time derivatives. Using $J_{n}=c_{0}\Lambda^{n0}$ and integrating by parts, equation (25) gives a local effective Lagrangian
$\mathcal{L}_{\rm eff}=L_{0}+\sum_{n\neq 0}\frac{J_{n}^{*}J_{n}}{\Delta_{n}}+\mathcal{O}(\partial_{t}).$ (27)
By substituting $J_{n}=c_{0}\dot{\bm{q}}_{c}\cdot\bm{A}_{[n0]}$ into the above equation and neglecting the Berry phase corrections to the band gap, i.e., $|\Lambda_{[nn]}-\Lambda_{[00]}|\ll|\omega_{[n0]}|$ and $\Delta_{n}\simeq\omega_{[n0]}$, we obtain the effective Lagrangian as shown in equation (6).
V Acknowledgments
@@
Y.R. thanks Qian Niu, Yang Gao, Di Xiao, Wenqin Chen and Jie Wang for useful discussions.